AN electric heater takes 4A when operated on a 250v supply. what is the cost of electricity consumed at 10K per KWh when the heater is used for 5 hours
1
Expert's answer
2020-06-18T10:57:51-0400

Let us calculate the power of the heater. It can be calculated as "P=UI = 250\,\mathrm{V}\cdot 4\,\mathrm{A} = 1000\,\mathrm{W} = 1\,\mathrm{kW}."

Therefore, if the heater is used for 5 hours, its energy will be "5\,\mathrm{h}\cdot 1\,\mathrm{kW} = 5\,\mathrm{kW\cdot h}" . So the cost will be "10\,\mathrm{K/KWh} \cdot 5\,\mathrm{kWh} = 50\,\mathrm{K}."